OpenStudy (anonymous):

PLEASE HELP!!!!!!! ;((( simplify: √-10 (+11+√-5)

OpenStudy (anonymous):

\[\sqrt{-10}(11+\sqrt{-5})\]?

OpenStudy (anonymous):

\[11\sqrt{10}i-\sqrt{50}\]

OpenStudy (anonymous):

or if you prefer \[-5\sqrt{2}+11\sqrt{10}i\]

OpenStudy (anonymous):

wouldnt it be -5√2 + 11i√10 @satellite73

OpenStudy (anonymous):

\[11\sqrt{10}i\] is the same as \(11i\sqrt{10}\) notice that the \(i\) is not inside the radical standard form of a complex number is \(a+bi\) where \(a\) and \(b\) are real numbers, so standard form would be \(-5\sqrt{2}+11\sqrt{10}i\) but they are the same